# Automated Layout Design and Control of Robust Cooperative Grasped-Load Aerial Transportation Systems

### v2

@@ Abstract
We present a novel approach to cooperative aerial transportation through a team of quadcopters, using optimal control theory and a hierarchical control strategy. We assume the quadcopters are connected to the payload through rigid attachments, essentially transforming the whole system into a larger flying object with “thrust modules” at the attachment locations of the quadcopters. We investigate the optimal arrangement of the thrust modules around the payload, so that the resulting system is robust to disturbances. We choose the $\mathcal{H}_{2}$ norm as a metric of robustness, and propose an iterative optimization routine to compute the optimal layout of the vehicles around the object. We experimentally validate our approach using four quadcopters and comparing the disturbance rejection performances achieved by two different layouts (the optimal one and a sub-optimal one), and observe that the results match our predictions.
@@ I INTRODUCTION
The growing demand for effective and autonomous transportation in applications such as logistics, search and rescue operations, and load lifting has increased interests in efficient aerial transportation solutions [1]. Being able to apply aerial robots to logistics and transportation would not only reduce costs and enhance efficiency, but also allow to reduce the burden on ground infrastructures. The application of UAVs to these fields is challenging, mainly due to their payload and battery life limitations. We address the problem of using multiple quadcopters for payload lifting and transportation. A lot of work has been carried out on design, control and path planning for single quadcopter transportation systems, some examples of which are [2, 3, 4, 5, 6, 7]. However, practical limitations on dimensions, vehicle complexity, load capacity, and costs limit the application of such technologies.
@@
The widespread use and cost-effectiveness of smaller vehicles, such as quadcopters, have made them the preferred choice for practical applications [8]. However, using multiple smaller vehicles in a cooperative manner, along with potential payload interactions (e.g. aerodynamic interactions, vibrations), also introduces greater complexity in terms of control and trajectory planning for cooperative aerial transportation systems [9, 10]. Connecting the vehicles to the payload through tethers is a popular choice, and allows to distance the agents from the payload. The dynamics and control of tethered systems have been extensively investigated and yielded interesting results, e.g. [11]. Other examples are [12], which employed an interconnected structure to enhance the stability of the tethered system, or [13], which implemented a decentralized control scheme not relying on assumptions about the payload’s state. [14] presented a cooperative control approach which showed significant disturbance rejection performances.
@@
Figure 1: Picture of four quadcopters cooperatively carrying a single panel payload.
Compared to the challenges of managing tethered or moving payloads, which can be complex due to their internal dynamics and limited maneuverability, rigid attachments are often the preferred option for transportation purposes. Some work has been carried out also along this direction. [15] studied the control for cooperative payload stabilization with a team of quadcopters equipped with grippers. [16] focused on cooperative grasped transportation control using onboard visual and inertial sensing, while [17] implemented an adaptive control framework based on a Kalman Filter for systems with uncertainty about quadcopter location and payload parameters. [18] explores the possibility of transporting flexible payloads, taking into account internal deformations and dynamics and aiming to minimize deformations.
Our research addresses the synthesis of rigidly attached transportation systems. The core concept is to address both the design layout and the optimal control of the combined quadcopter-payload system as one unified problem. We do this by formulating and solving an optimization problem derived from robust control theory for a first order model of the system. In particular, $\mathcal{H}_{2}$ control proved to be a promising approach to maximize disturbance rejection performances [19]. The main contribution of this research is an automated design tool to determine the optimal thrust module arrangement for payload stability and robustness. Experimentally validated with four quadcopters, our method confirms the effectiveness of the optimal layout in rejecting disturbances.
The paper is structured as follows. In Section II we describe the mathematical framework employed to compute the optimal placement locations and the linear feedback law to control the system. We then describe the control infrastructure adopted in our hardware setup. In Section III we show some example outputs of our optimization routine applied to different payload shapes, masses and number of quadcopters used for transportation. In Section IV we introduce the experimental setup and the flight tests which were carried out to validate the theoretical predictions. Finally, in Section V we discuss the results, relevance of the work and future developments.
II MATERIALS AND METHODS
The problem we solve is twofold. We first optimize the placement of the quadcopters on the object to be transported, so as to maximize the system’s disturbance rejection performance. This involves solving an optimization problem where the cost function is derived from $\mathcal{H}_{2}$ control theory, and the decision variables are the quadcopters’ attachment locations. We then address the constrained quadcopters dynamics through a hierarchical control strategy of the integrated quadcopter-payload system. The motion of the system is controlled using the gain obtained from the same $\mathcal{H}_{2}$ control used for layout optimization. Each quadcopter (thrust module) is sent thrust commands. An example picture of a team of quadcopters carrying a payload with rigid attachments is shown in Fig. 1.
@@ II-A Modelling
We make the two following key assumptions about our system.
Assumption 1
The payload is characterized by an orthogonal extrusion geometry.
Assumption 2
The payload inertia is fully known.
Assumption 1 simplifies the following mathematical derivation, but does not limit significantly the space of possibilities: indeed many objects, such as delivery packages, boxes, and pallets are all included in this category. In such case, we investigate the quadcopters’ placement on the mid plane of the extrusion (Fig. 2a)), reducing the size of the optimization problem formulated. The relaxation of this assumption will be a future extension. Assumption 2 simplifies the control infrastructure, and can also be relaxed in the future.
With reference to Fig. 2b), the placement variables, are defined as the angles $\boldsymbol{\theta}=[\theta_{i},...,\theta_{N}]^{T}$ describing the attachments locations along the curve $\Gamma$ with respect to the centroid of the payload shape. We highlight that this is just a geometric parametrization choice, and no assumption is made on the overall center of gravity of the system. The attachment is assumed to be obtained through a rigid rod, which allows to keep distance between payload and thrust modules.
The quadcopter-payload ensemble is itself a rigid body, and can therefore be described by a state vector $\mathbf{x}=[\mathbf{p}^{T},\dot{\mathbf{p}}^{T},\boldsymbol{\gamma}^{T},\boldsymbol{\omega}^{T}]^{T}\in\mathbb{R}^{12}$, where $\mathbf{p}$ and $\dot{\mathbf{p}}$ are the position and velocity of the centroid of the object with respect to a global world reference frame, $\boldsymbol{\gamma}$ is the Roll-Pitch-Yaw triplet defining the orientation of the local frame, and $\boldsymbol{\omega}$ the angular velocity with respect to the local body frame. Tracking the motion of the centroid of the payload rather than the center of mass of the system (which in general do not coincide) is a design choice due to the geometric relevance of the former. As such, our methods would not be impacted by a different definition of the state vector. Each quadcopter is seen as a thrust module providing four inputs to the system. Therefore, if $N$ quadcopters are deployed, the thrust input vector is $4N$-dimensional, i.e. $\mathbf{u}\in\mathbb{R}^{4N}$. Due to unavoidable hardware limitations, we assume that each thrust component is limited to the interval $[u_{l},u_{h}]$. It is possible to linearize the rigid body dynamics model about hover configuration, and decompose the inputs in $\mathbf{u}=\bar{\mathbf{u}}+\mathbf{u^{\prime}}$, in which $\bar{\mathbf{u}}$ are the feedforward thrusts and $\mathbf{u}^{\prime}$ are the first order components of the linearized model. We do the same for the state vector $\mathbf{x}=\bar{\mathbf{x}}+\mathbf{x}^{\prime}$. We model the disturbances applied to the system as a random variable $\mathbf{d}=[\mathbf{F}_{d}^{T},\mathbf{T}_{d}^{T}]^{T}$, in which $\mathbf{F}_{d}\in\mathbb{R}^{3}$ and $\mathbf{T}_{d}\in\mathbb{R}^{3}$ represent white noise force and a torque applied to the origin of the local body reference frame. The problem is formulated as the minimization of the probability of infeasible inputs when the system is subject to the defined random disturbance vector $\mathbf{d}\in\mathbb{R}^{6}$.
@@
Figure 2: a) Extrusion payload geometry and representation of the mid-plane. Its intersection with the side faces is the curve $\Gamma$. b) Schematics of four quadcopters attached around the payload along $\Gamma$. Representation of the local body reference frame centered at the centroid of $\Gamma$, placement variables defined by the angles $\boldsymbol{\theta_{i}}$ ($i=1,...,4$), and disturbance force $F_{d}$ and torque $T_{d}$.
@@
We analyze the integrated quadcopter-payload system as a whole. We can write its continuous-time linearized dynamics as:
@@
The dynamics matrix $A\in\mathbb{R}^{12\times 12}$ is obtined by linearization of the rigid body dynamics equations about hovering and does not depend on the placement variables $\boldsymbol{\theta}$. The matrices $B(\boldsymbol{\theta})\in\mathbb{R}^{12\times 4N}$ and $B_{d}(\boldsymbol{\theta})\in\mathbb{R}^{12\times 6}$ depend on $\boldsymbol{\theta}$ either directly or through the inertia of the system (which itself is a function of $\boldsymbol{\theta}$).
II-B Layout Optimization
With reference to the $\mathcal{H}_{2}$ optimal control literature [20], we define an additional auxiliary variable:
@@ II-B Layout Optimization
The dimension of $\mathbf{z}$ can vary, and mainly depends on the desired control authority along the directions of interest in the state and input spaces. The goal is to find the feedback gain $K^{*}$ such that
$\mathbf{u}=\bar{\mathbf{u}}-K^{*}(\boldsymbol{\theta})\mathbf{x}^{\prime},$ (3)
@@
$K^{*}(\boldsymbol{\theta})=(D^{T}D)^{-1}B(\boldsymbol{\theta})^{T}S_{1}(\boldsymbol{\theta})$ (4)
where $S_{1}(\boldsymbol{\theta})$ is the solution of the algebraic Riccati equation:
$A^{T}S_{1}+S_{1}A-S_{1}B(D^{T}D)^{-1}B^{T}S_{1}+C^{T}C=0$ (5)
@@
in which we omitted the dependency on $\boldsymbol{\theta}$ for clarity of writing. The feedback gain does not depend on the matrix $B_{d}$, which represents how the disturbance affects the dynamics. The optimal cost value turns out to be:
$J^{*}=\text{tr}(B_{d}^{T}S_{1}B_{d}).$ (6)
@@
The optimal cost value does indeed depend on the matrix $B_{d}$. This is a valid cost function which can be used to optimize the quadcopters’ attachment locations. However, it does not guarantee that the inputs generated by eq. 3 are feasible. Therefore, we make an additional step: we compute the covariance $S_{2}(\boldsymbol{\theta})$ of the state vector in a condition in which the optimal feedback is active and the system is subject to white noise disturbance. This can be computed using the optimal continuous-time state observer theory, which leads to the equation:
$A_{f}S_{2}+S_{2}A_{f}^{T}+B_{d}B_{d}^{T}=0$ (7)
@@
where we defined $A_{f}(\boldsymbol{\theta})=A-B(\boldsymbol{\theta})K^{*}(\boldsymbol{\theta})$ as the dynamics matrix with linear feedback. Note that the assumption of white noise disturbance makes the process ergodic, and therefore the state sample mean is equivalent to the temporal mean. The matrix $S_{2}(\boldsymbol{\theta})$ is then an approximation of the covariance of the first order state linearization $\mathbf{x}^{\prime}$ when subject to white noise disturbance. At this point we have the feed-forward thrust values, which are computed from the hovering equilibrium condition, and the covariance matrix of the inputs, obtained through the feedback law $K^{*}(\boldsymbol{\theta})^{T}S_{2}(\boldsymbol{\theta})K^{*}(\boldsymbol{\theta})$. The goal is to maximize the probability of feasible inputs. If we set the element-wise thrust lower bound to be $\mathbf{u}_{l}=u_{l}\mathbf{1}_{4N}$ and the upper bound to be $\mathbf{u}_{h}=u_{h}\mathbf{1}_{4N}$, where $\mathbf{1}_{4N}$ indicates the $4N$-dimensional vector of ones. We can then write the optimization problem we are trying to solve as:
$\max_{\boldsymbol{\theta}}F(\mathbf{u}_{h})-F(\mathbf{u}_{l})$ (8)
@@
Figure 3: Two dimensional example of minimum Mahalanobis distance to input bounds. In this case, the random variables are $u_{1},u_{2}\in[u_{l},u_{h}]$. The ellipses depicted, centered at the mean $\bar{\mathbf{u}}$, represent the level curves of the Mahalanobis distance and are obtained as $(\mathbf{u}-\bar{\mathbf{u}})^{T}\Sigma_{\mathbf{u}}^{-1}(\mathbf{u}-\bar{\mathbf{u}})=q$ ($q$ being a positive scalar parameter). The “point of minimum distance” sought is given by the tangency between the largest ellipsoid contained in the bounding box and the corresponding bounding hyperplane.
@@
where $F(\cdot)$ is the cumulative distribution over thrust inputs, and its mean and covariance depend on $\boldsymbol{\theta}$. It is possible to approximate this cumulative distribution through sampling, but the sampling approach would lead to a noisy cost function for our optimization problem, whose variance decreases linearly with the number of samples used. An additional limitation of sampling is that it needs a distribution to draw samples from, and this would have to be assumed. In addition, sampling-based methods are computationally expensive and lead to long solution times. Therefore, we choose to leverage the concept of minimum Mahalanobis distance [21] of the mean (feedforward thrusts) to the thrust bounds. The Mahalanobis distance of a point $\mathbf{u}$ to a distribution with mean $\bar{\mathbf{u}}$ and covariance $\Sigma_{\mathbf{u}}$ is defined as:
@@
We can extend the concept of Mahalanobis distance to a hyperplane as the minimum Mahalanobis distance of a point of the hyperplane from the distribution. A two dimensional example is shown in Fig. 3. Since in our case we have a lower and an upper bound for each input, and four thrust inputs per quadcopter, there are $8N$ different bounding hyperplanes to consider in the input space. Therefore, going back to the goal of maximizing the probability of feasible inputs, we aim to maximize the Mahalanobis distance of the mean (feedforward thrusts) to the thrust bounds. The optimization problem can be formulated as:
$\max_{\boldsymbol{\theta}}\biggl{\{}\,\,\min_{k}$ $\left[\min_{u_{a}\in\{u_{l},u_{h}\}}\biggl{(}\,\,\min_{\mathbf{u}}d_{M}\Bigl{(}\mathbf{u};\bar{\mathbf{u}}(\boldsymbol{\theta}),\Sigma_{\mathbf{u}}(\boldsymbol{\theta})\Bigl{)}\,\,\biggl{)}\,\,\right]\,\,\biggl{\}}$ (10)
@@
$\,\,\,\,\,\,\,\,\,u^{(k)}=u_{a}$
In which with the superscript $u^{(k)}$ we refer to the $k$-th component of the vector $\mathbf{u}$. We are then dealing with a nested optimization problem. Trying to unpack what the routine is about: we are looking for the value $\boldsymbol{\theta}^{*}$ which maximizes the minimum Mahalanobis distance of the thrust bounds from the feedforward thrust. This minimum distance is computed by evaluating it on each bounding hyperplane separately, hence the minimization over $k$ (component-wise) and over $u_{a}\in\{u_{l},u_{h}\}$, where the subscript $a$ stands for “active”. The inner minimization over $\mathbf{u}$ is necessary because, on each bounding hyperplane, we are interested in finding the point with minimum distance from $\bar{\mathbf{u}}$. The optimization routine is also described in Algorithm 1. On the computational cost level, there are no particularly expensive operations. Solving Riccati equations is relatively fast, and minimizing the Mahalanobis distance from a hyperplane is a Quadratic Program (QP). For each update of $\boldsymbol{\theta}$ in the optimization routine, $8N$ QPs are solved. The solver used for the outer optimization loop is the Nelder-Mead simplex method discussed in [22] and [23].
@@
     Compute $S_{1}(\boldsymbol{\theta})$ and $K^{*}(\boldsymbol{\theta})$ from eq. 5
     Compute $S_{2}(\boldsymbol{\theta})$ from eq. 7
@@
              Solve QP $d^{\prime}\leftarrow\min_{\mathbf{u}}d_{M}(\mathbf{u};\bar{\mathbf{u}},\Sigma_{\mathbf{u}})$ as in eq.
                            10 s.t. $\mathbf{u}_{l}\leq\mathbf{u}\leq\mathbf{u}_{h}$, $u^{(k)}=u_{a}$
@@
     Update $\boldsymbol{\theta}$ (Nelder–Mead simplex algorithm)

### v3

# Automated Layout and Control Co-Design of Robust Multi-UAV Transportation Systems

@@ Abstract
The joint optimization of physical parameters and controllers in robotic systems is challenging. This is due to the difficulties of predicting the effect that changes in physical parameters have on final performances. At the same time, physical and morphological modifications can improve robot capabilities, perhaps completely unlocking new skills and tasks. We present a novel approach to co-optimize the physical layout and the control of a cooperative aerial transportation system. The goal is to achieve the most precise and robust flight when carrying a payload. We assume the agents are connected to the payload through rigid attachments, essentially transforming the whole system into a larger flying object with “thrust modules” at the attachment locations of the quadcopters. We investigate the optimal arrangement of the thrust modules around the payload, so that the resulting system achieves the best disturbance rejection capabilities. We propose a novel metric of robustness inspired by $\mathcal{H}_{2}$ control, and propose an algorithm to optimize the layout of the vehicles around the object and their controller altogether. We experimentally validate the effectiveness of our approach using fleets of three and four quadcopters and payloads of diverse shapes.
@@ I Introduction
Autonomous aerial transportation solutions have been proven increasingly essential in applications such as construction, logistics, and load lifting [1]. Being able to scalably and reliably apply aerial robots to these settings would not only drastically reduce costs and enhance time and energy efficiency, but also reduce the need of ground infrastructures. The deployment at scale of Uncrewed Aerial Vehicles (UAVs) to these scenarios is challenging, mainly due to their load capacity and robustness limitations, especially significant when carrying a payload. We address the problem of synthesizing a maximally robust multi-UAV system for collaborative payload lifting and transportation.
A lot of work has been carried out on design, control and path planning for single quadcopter transportation systems, some examples of which are [2, 3, 4, 5]. However, practical limitations on dimensions, vehicle complexity, load capacity, and costs limit the application of such technologies. The widespread use and cost-effectiveness of smaller vehicles, such as quadcopters, have made them the preferred choice for practical applications [6]. Their use in a collaborative fashion has been proposed, for instance, for construction applications, but in simplified settings in which each agent individually carries a small load [7, 8].
@@
Using multiple small vehicles in a cooperative manner allows to increase the payload, but also introduces greater complexity in terms of control, due to e.g. aerodynamic interactions and vibrations, and trajectory planning [9, 10]. Connecting the vehicles to the payload through tethers is a popular choice [11], and allows to distance the agents from the payload. The dynamics and control of tethered systems have been extensively investigated, e.g. in [12, 13]. Other examples are [14], which employed an interconnected structure to enhance the system stability, [15], which proposed a decentralized approach to the system’s control, and [16], where improved disturbance rejection capabilities were achieved.
Compared to the challenges of managing tethered or moving payloads, which can be complex due to their internal dynamics and limited maneuverability, rigid attachments are often the preferred option for transportation purposes. Researchers have investigated hardware aspects, such as adding grippers to the vehicles [17], and control aspects, such as limited sensing capabilities [18], adaptive control frameworks [19], and compensation of internal payload vibrations [20].
@@
Figure 1: Four quadcopters cooperatively carrying a single panel payload.
The use of multiple thrust modules [21, 22, 23] introduces additional parameters to the system, which can be leveraged to enhance performances. The physical configuration of the lifting agents around a payload is particularly interesting, because it can drastically influence the robustness during closed-loop flight. Selecting the best layout, however, can be complex, due to the intricate dependencies of flight performance metrics on physical parameters (which affect the system dynamics, often in a nonlinear fashion). In previous works, such as [17, 21, 19], this is done through simple heuristics, or by intuition, and compensating inaccuracies through adaptive control approaches, leading to suboptimal flying systems.
An area addressing such joint control-hardware optimization problems to achieve superior performances is control co-design [24]. These methodologies have been successfully applied to a number of robotics applications, such as hands [25], legged robots [26], multicopters [27, 28], and winged drones [29, 30]. The output of these methods is typically a control policy paired with a set of physical system parameters which jointly maximize the desired objective.
Our research applies a co-design approach to the synthesis of collaborative aerial transportation systems. The core idea is to jointly compute the layout and the control of the combined quadcopter-payload system as a single unified problem, maximizing robustness. Our goal is to automate the design process and synthesize systems with high performances by tuning its physical parameters. We do this by formulating and solving an optimization problem inspired by robust control theory for a first-order system model, method that has previously been used to capture disturbance rejection performances of a quadcopter [31]. The main contribution of this paper is an efficient co-design tool to optimally arrange thrust modules around a payload to maximize robustness during flight, and consists of two major parts:
  • A novel cost function capturing the disturbance response and the system’s input constraints using state-of-the-art methodologies;
  • A computationally efficient procedure to determine the optimal thrust module arrangement based on the proposed cost function.
II Methodology
The co-design problem we solve is intrinsically twofold, as it involves optimizing over a physical configuration as well as over control parameters. The question we ask is: what is the physical placement of $N$ quadcopters around a payload, and associated controller, such that, the overall flying system is the most robust possible to disturbances? We cast the problem as an optimization problem where the cost function is inspired by $\mathcal{H}_{2}$ control theory, and the decision variables are the quadcopters’ attachment locations. In the inner loop of the optimization, a Linear Quadratic Regulator (or LQR, i.e. the $\mathcal{H}_{2}$-optimal controller) is associated to candidate layouts for evaluation. This choice, in fact, allows a straightforward computation of the control (i.e. a linear feedback matrix), and also an analytic evaluation of the $\mathcal{H}_{2}$-inspired cost function. More advanced control strategies could be used, at the expense of a more involved policy optimization, and a more computationally expensive evaluation of the cost function (using for example Monte Carlo methods). We make the assumption that a reliable system for state estimation is available.
@@ II-A Modelling
We assume that the payload is characterized by a right prism geometry, whose mass properties and shape are known. These assumptions simplify the following mathematical derivation, but do not limit the space of application. Indeed, many common payloads, such as packages, boxes, pallets all satisfy these assumptions. We investigate the quadcopters’ placement on the payload mid-plane (Fig. 2a), reducing the amount of decision variables in the optimization problem formulated. With reference to Fig. 2b, the placement variables, are defined as the angles $\boldsymbol{\theta}=[\theta_{i},...,\theta_{N}]^{T}$ describing the attachments locations along the curve $\Gamma$ with respect to the centroid of the payload shape. The attachment is assumed to be obtained through a rigid rod, which allows to keep distance between payload and thrust modules.
The quadcopter-payload ensemble is itself a rigid body, and can therefore be described by a state vector $\mathbf{x}=[\mathbf{p}^{T},\dot{\mathbf{p}}^{T},\boldsymbol{\gamma}^{T},\boldsymbol{\omega}^{T}]^{T}\in\mathbb{R}^{12}$, where $\mathbf{p}$ and $\dot{\mathbf{p}}$ are the position and velocity of the centroid of the object with respect to a global world reference frame, $\boldsymbol{\gamma}$ is the Roll-Pitch-Yaw triplet defining the orientation of the local frame (this is chosen here for ease of exposition), and $\boldsymbol{\omega}$ the angular velocity with respect to the local body frame.
Each quadcopter is seen as a thrust module providing four inputs to the system. Therefore, if $N$ quadcopters are deployed, the thrust input vector is $4N$-dimensional, i.e. $\mathbf{u}\in\mathbb{R}^{4N}$. We assume that each thrust component is limited to the interval $[u_{l},u_{h}]\subset\mathbb{R}$.
For the derivation of the dynamics equations in vector form, we follow [32]. The motion is described by
$\begin{cases}m\ddot{\mathbf{p}}=m\mathbf{g}+R\,\mathbf{e}_{3}T\\ J\dot{\boldsymbol{\omega}}+\boldsymbol{\omega}\times J\boldsymbol{\omega}=\boldsymbol{\tau}\end{cases},$ (1)
where $m$ and $J$ are the total mass and inertia matrix of the system, $\mathbf{g}=[0,0,-9.81]^{T}$ is the gravity acceleration vector, $R$ is the rotation matrix from (3) in [32] and obtained from the triplet $\boldsymbol{\gamma}$, $\mathbf{e}_{3}$ is the third vector of the standard Euclidean basis, $T=\sum_{i=1}^{4N}u_{i}$ is the total thrust produced by all quadcopters, $\mathbf{1}_{4N}$ is the $4N$-dimensional vector of ones, $\boldsymbol{\tau}$ is the total torque vector in the local body frame. The vector $\boldsymbol{\tau}$ is obtained by computing the thrust and torques developed by each quadcopter (from (4) and (5) in [32]) and translating them to the local body reference frame through the known relative position vectors of the attachment points.
We linearize the rigid body dynamics model about hover configuration and decompose the inputs in $\mathbf{u}=\bar{\mathbf{u}}+\mathbf{u^{\prime}}$, in which $\bar{\mathbf{u}}$ are the feedforward thrusts and $\mathbf{u}^{\prime}$ are the first order components of the linearized model. We do the same for the state vector $\mathbf{x}=\bar{\mathbf{x}}+\mathbf{x}^{\prime}$. We model the disturbances applied to the system as random Gaussian noise in the form $\mathbf{d}=[\mathbf{f}_{d}^{T},\mathbf{t}_{d}^{T}]^{T}$, in which $\mathbf{f}_{d}\in\mathbb{R}^{3}$ and $\mathbf{t}_{d}\in\mathbb{R}^{3}$ represent disturbance force and a torque applied to the origin of the local body reference frame (shown in Fig. 2b).
@@
Figure 2: a) Extrusion payload geometry and representation of the mid-plane. Its intersection with the side faces is the curve $\Gamma$. b) Schematics of four quadcopters attached around the payload along $\Gamma$. Representation of the local body reference frame centered at the centroid of $\Gamma$, placement variables defined by the angles $\boldsymbol{\theta_{i}}$ ($i=1,...,4$), and disturbance force $\mathbf{f}_{d}$ and torque $\mathbf{t}_{d}$.
@@
We can write the continuous-time linearized dynamics of the quadcopter-payload system (as a whole) as:
@@
The dynamics matrix $A\in\mathbb{R}^{12\times 12}$ is obtained by taking the Jacobian of (1) around hovering conditions with respect to the state vector $\mathbf{x}$. $A$ does not depend on the placement variables $\boldsymbol{\theta}$. The matrices $B(\boldsymbol{\theta})\in\mathbb{R}^{12\times 4N}$ and $B_{d}(\boldsymbol{\theta})\in\mathbb{R}^{12\times 6}$ are obtained by taking the Jacobian of (1) with respect to the input $\mathbf{u}$ and disturbance $\mathbf{d}$ vectors, respectively. These matrices depend on $\boldsymbol{\theta}$ directly or through the inertia of the system (which itself is a function of $\boldsymbol{\theta}$). It is also important to note that the error due to the linearization of the dynamics equations does not depend on $\boldsymbol{\theta}$, and the linear approximation is reasonably accurate for the applications of interest in this paper.
II-B Control and layout co-design
In this section, we present our co-design approach. First, we show how to derive the optimal $\mathcal{H}_{2}$ controller for the system of interest, and then how to use this approach to maximize robustness with respect to the layout variables. An alternative approach could be the $\mathcal{H}_{\infty}$ control [33], which by design is optimal for the worst-case disturbance. However, we find that $\mathcal{H}_{2}$ control is a better choice in our case. In fact, the worst-case disturbance for the systems of interest do not match the ones that are actually seen at deployment, which are better approximated as white noise.
With reference to the $\mathcal{H}_{2}$ optimal control literature [33], we define an additional auxiliary variable
@@ II-B Control and layout co-design
The goal is to find the feedback gain $K^{*}\in\mathbb{R}^{4N\times 12}$ (dependent on $\boldsymbol{\theta}$) such that
$\mathbf{u}^{\prime}=-K^{*}(\boldsymbol{\theta})\mathbf{x}^{\prime},$ (4)
@@
It is a known result that the optimal linear feedback controller minimizing the cost function $J$ is obtained as
$K^{*}(\boldsymbol{\theta})=(D^{T}D)^{-1}B(\boldsymbol{\theta})^{T}S_{1}(\boldsymbol{\theta}),$ (5)
where $S_{1}(\boldsymbol{\theta})$ is the solution of the algebraic Riccati equation
$A^{T}S_{1}+S_{1}A-S_{1}B(D^{T}D)^{-1}B^{T}S_{1}+C^{T}C=0.$ (6)
@@
We omitted the dependency on $\boldsymbol{\theta}$ in (6) for clarity of writing. The feedback gain does not depend on the matrix $B_{d}$, which represents how the disturbance affects the dynamics. The optimal cost value is
$J^{*}=\text{trace}(B_{d}^{T}S_{1}B_{d}),$ (7)
@@
which instead does depend on the matrix $B_{d}$. This can be used to optimize the quadcopters’ attachment locations. However, this formulation only considers feedback and ignores feedforward terms, and does not guarantee that the inputs generated by (4) are feasible. Therefore, we do an additional step: we compute the closed-loop state covariance $S_{2}(\boldsymbol{\theta})$ of the system controlled with the optimal feedback defined in (4), subject to the same white noise disturbance $\mathbf{d}$. This can be computed using the optimal continuous-time state observer theory, which leads to the equation:
$A_{f}S_{2}+S_{2}A_{f}^{T}+B_{d}B_{d}^{T}=0,$ (8)
@@
where we defined $A_{f}(\boldsymbol{\theta})=A-B(\boldsymbol{\theta})K^{*}(\boldsymbol{\theta})$ as the closed-loop dynamics matrix using the linear feedback of (4). Note that the Gaussian noise disturbance makes the process ergodic, and therefore the state sample mean is equivalent to its temporal mean. The matrix $S_{2}(\boldsymbol{\theta})$ is then an approximation of the covariance of the first order state linearization $\mathbf{x}^{\prime}$ when subject to white noise disturbance.
At this point, we have the feed-forward thrust values, which are computed from the hovering equilibrium condition, and the covariance matrix of the inputs $\Sigma_{\mathbf{u}}=K^{*}(\boldsymbol{\theta})^{T}S_{2}(\boldsymbol{\theta})K^{*}(\boldsymbol{\theta})$, obtained through the feedback law. The goal is to maximize the probability of feasible inputs. We set the element-wise thrust lower saturation to be $\mathbf{u}_{l}=u_{l}\mathbf{1}_{4N}$ and the upper saturation to be $\mathbf{u}_{h}=u_{h}\mathbf{1}_{4N}$, where $\mathbf{1}_{4N}$ indicates the $4N$-dimensional vector of ones. We can then write the optimization problem we are trying to solve as:
$\max_{\boldsymbol{\theta}}\bigl{\{}F(\mathbf{u}_{h})-F(\mathbf{u}_{l})\bigr{\}}$ (9)
@@
Figure 3: Two dimensional example of minimum Mahalanobis distance to input saturations. In this case, the random variables are $u_{1},u_{2}\in[u_{l},u_{h}]$. The ellipses depicted, centered at the mean $\bar{\mathbf{u}}$, represent the level curves of the Mahalanobis distance and are obtained as $(\mathbf{u}-\bar{\mathbf{u}})^{T}\Sigma_{\mathbf{u}}^{-1}(\mathbf{u}-\bar{\mathbf{u}})=q$ ($q$ being a positive scalar parameter). The “point of minimum distance” sought is given by the tangency between the largest ellipsoid contained in the bounding box and the corresponding saturation hyperplane.
@@
where $F(\cdot)$ is the cumulative distribution over thrust inputs, and its mean and covariance depend on $\boldsymbol{\theta}$. It is possible to approximate this cumulative distribution through sampling, but the sampling approach would lead to a noisy cost function for our optimization problem, whose variance decreases linearly with the number of samples used. In addition, sampling-based methods are computationally expensive and lead to long solution times. Therefore, we choose to leverage the concept of Mahalanobis distance [34] to quantify the margin between mean (feedforward thrusts) and the thrust saturations. The Mahalanobis distance of a point $\mathbf{u}$ to a distribution with mean $\bar{\mathbf{u}}$ and covariance $\Sigma_{\mathbf{u}}$ is defined as:
@@
We extend the concept of Mahalanobis distance of a hyperplane from a distribution as the minimum Mahalanobis distance of a point of the hyperplane from that distribution. A two dimensional example is shown in Fig. 3. Since in our case we have a lower and an upper saturation ($u_{l}$ and $u_{h}$ respectively) for each input component, and four thrust inputs per quadcopter, there are $8N$ different hyperplanes to consider in the input space. Therefore, going back to the goal of maximizing the probability of feasible inputs, we aim to maximize the Mahalanobis distance of the thrust saturation hyperplanes from the input distribution. The optimization problem can be formulated as:
$\max_{\boldsymbol{\theta}}\biggl{\{}\,\,\min_{\mathbf{u}}$ $\biggl{[}\,\,d_{M}\Bigl{(}\mathbf{u};\bar{\mathbf{u}}(\boldsymbol{\theta}),\Sigma_{\mathbf{u}}(\boldsymbol{\theta})\Bigl{)}\,\,\biggl{]}\,\,\biggl{\}}$ (11)
@@
We are looking for the value $\boldsymbol{\theta}^{*}$ which maximizes the minimum Mahalanobis distance of the thrust saturation hyperplanes from the feedforward thrust. This minimum distance is computed by evaluating it on each saturation hyperplane separately, hence this minimization step is component-wise and performed over both saturations $\{u_{l},u_{h}\}$. The inner minimization over $\mathbf{u}$ is necessary because, on each hyperplane, we are interested in finding the point with minimum distance from $\bar{\mathbf{u}}$. The optimization routine is also described in Algorithm 1. With the superscript expression $u^{(k)}$ we refer to the $k$-th component of the vector $\mathbf{u}$, while the expression $u_{a}\in\{u_{l},u_{h}\}$ is used to denote which of the two thrust saturation constraints is active at a given stage of the optimization.
On the computational cost level, there are no particularly expensive operations. Solving Riccati equations is relatively fast, and minimizing the Mahalanobis distance from a hyperplane is a Quadratic Program (QP). For each update of $\boldsymbol{\theta}$ in the optimization routine, $8N$ QPs are solved. The solver used for the outer optimization loop is the Nelder-Mead simplex method, details of which can be found in [35] and [36]. In general, this optimization problem is non-convex, even if the payload shape is convex. A trivial example is the invariance to permutation of the $\boldsymbol{\theta}$ vector, which alters the optimizer but not the optimum. There are also cases in which local minima exist. To make sure the minimum is found in practice, we run the algorithm multiple times with different initial guesses.
An important observation to make is that our algorithm deals with feedback and feedforward in a joint fashion, maximizing the margin from actuator saturation when the system is in closed-loop. The optimal LQR controller is re-computed at each iteration and is thus a byproduct of our co-design tool.
@@
     Compute $S_{1}(\boldsymbol{\theta})$ and $K^{*}(\boldsymbol{\theta})$ from (6)
     Compute $S_{2}(\boldsymbol{\theta})$ from (8)
@@
              Solve QP (as in (LABEL:eq:optimization)):
                                   $d^{\prime}\leftarrow\min_{\mathbf{u}}d_{M}(\mathbf{u};\bar{\mathbf{u}},\Sigma_{\mathbf{u}})$
                                                   s.t. $\mathbf{u}_{l}\leq\mathbf{u}\leq\mathbf{u}_{h}$, $u^{(k)}=u_{a}$
@@
     Update $\boldsymbol{\theta}$ (Nelder-Mead simplex algorithm)